Which of the following solutes will produce the largest increase in boiling point upon addition to 1.0 kg of water: (1.0 mole of Co(NO3)2, 2.0 moles of KCl or 3.0 moles of C2H5OH)?
1
Expert's answer
2021-08-10T06:53:24-0400

1 kg water = 1L = 1000

Molar mass of CO = 28.01

Molar mass of KCl = 74.55

Molar mass of C2H50H = 46.07

= 74.55×2/1000

= 0.1491

46.07/1000 = 0.13821

28.01*3/1000 = 0.02801


Hence, 2.0 moles of KCl